OpenStudy (anonymous):

The width of a rectangular poster is half its length. The area of the poster is 128 square feet. Let L = length.

OpenStudy (anonymous):

Let Length =L, Width = W, Area =A. We know that: \[A=WL\] and we are told that \[W=\frac{L}{2}\] so by substitution \[128=\frac{L}{2}L=\frac{L^{2}}{2}\]\[256=L^{2}\]\[L=\sqrt{256}\]\[L=16\]\[W=\frac{L}{2}=\frac{16}{2}=8\] Check\[128=LW=18*8=128\] CHECK